A population has a mean of 75 and a standard deviation of 17.4. A sample of 35 items is randomly selected from this population. What is the probability that the mean of this samples lies between 72 and 79? 


1
Expert's answer
2022-02-07T16:44:17-0500

Let "X=" the mean of the samples: "X\sim N(\mu, \sigma^2/n)"

Given "\mu=75, \sigma=17.4, n=35"


"P(72<X<79)=P(X<79)-P(X\le 72)"

"=P(X<\dfrac{79-75}{17.4/\sqrt{35}})-P(X\le \dfrac{72-75}{17.4/\sqrt{35}})"

"\approx P(X<1.360018)-P(X\le-1.020014)"

"\approx0.91309-0.15386=0.7592"